Can someone help me please. The problem is Complete the square. Then write the trinomial square in factored form. x^2 -2/5x
Answer by richwmiller(9144)

add (!/2*2/5)^2 to both sides
that equals (1/5)^2= 1/25
x^2-2x/5+1/25=1/25
factor 1/25
1/25 (25x^2-10x+1) = 1/25
multiply by 25
(25x^2-10x+1) = 1
(5x-1)^2=1


5x-1=1
5x=2
x=2/5
5x-1=-1
5x=0
x=0
x=0 x=2/5